Question1) identify the name of the compound represented on the spectrum graph?

Question2) For CNMR, indicate how many types of carbon atoms present and the identity of characteristics peaks?

Question3) For HNMR, indicate ho many types of carbon atoms present, the different identity of characteristics peaks, and explanations of it various splitting patterns?

Question4) Also for IR, indicate the identity of it characteristics peaks

Question5) Finally for MS, Identify the parent peak, the molecular weight and it implication, the most abundant peak, and identify the peak that indicate the presence of bromine or chlorine?


30 15 20 25 30 35 4045 50 3 6 70
Answer #1

The molecule has molecular ion peak at m/z = 58

Using rule 13,

58/13 = 4(6/13) ==> no of carbon = 4, no of hydrHydr = 4+6= 10

So, molecular formula = C4H10

Now, 13 showed two type of Carbon in molecule

And 1H NMR showed two peak in 9:1 which means 9 equivalent proton (3 X CH3) and one single proton (CH) is present in molecule.

So, the compound may be

(CH3)3CH = CH3--CH(CH3)--CH3

1) name of the compound 2-Methyl-propane

2)in 13C NMR two type of Carbon is present

at 30 ppm for CH3

at 50 ppm for CH

3)in 1H NMR two type of Hydrogen (I.e.two type Carbon atom) is present.

The peak at 1.0 ppm for CH3 = 9H, doublet (using n+1 rule, where n is no of Hydrogen)

The peak at 1.5 ppn for CH = 1H, decet.

4)in IR the peaks between 2900-3000 cm-1 is for CH stretching

Similarly, the peaks between 1200-1500 cm-1 for various bending and symmetric stretching of CH.

No functional group present in molecule.

5)The parent and most abundant peak in mass spectra at m/z =58

Use of this to identify the molecular formula as discussed in beginning.

Since, no Cl or Br Atom is present in molecule, so no peak in mass spectra.

Thank you